# Learning Interactive Real-World Simulators

### v2

@@ 11.2 Details of RL Policy Training
Stage 1 (Supervised Learning) Model Architecture The PaLI 3B model trained on Language-Table uses a Vision Transformer architecture G/14 [5] to process images, and the encoder-decoder architecture of UL2 language model [76] for encoding task descriptions and decoding tokens which can represent language, control actions, or other values of interest (described below). Objectives In the first stage of training, using a dataset of demonstrations, we finetune the pretrained PaLI 3B vision language model checkpoint [38] with the following tasks:
@@
  • Behavioral Cloning: Given observations and task instruction, predict the demonstration action. The continuous actions of the Language-Table domain are discretized into the form “+1 -5", and represented using extra tokens from the PaLI model’s token vocabulary. As an example, “+1 -5" is represented by the token sequence (<extra_id_65>, <extra_id_1>, <extra_id_66>, <extra_id_5>).
@@
We use learning rate 0.001, dropout rate 0.1, and batch size 128 to finetune the PaLI 3B model for 300k gradient steps with 1k warmup steps on both the simulated and real Language Table dataset similar to RT-2 brohan2023rt [32].
Stage 2 (RL Training) Reward Definition As mentioned above, during Stage 1, given an observation and goal, the PaLI model is finetuned to predict how many timesteps are left until the demonstration episode reaches a success state. Let us denote this function by $d(o,g)$. The reward we use during RL training is defined as $r(o_{t},a_{t},o_{t+1},g)=-[d(o_{t+1},g)-d(o_{t},g)]\cdot\mathcal{C}$, where $\mathcal{C}>0$ is a small constant used to stabilize training ($\mathcal{C}=5e-2$ in this work). Intuitively, this reward tracks if from timestep $t$ to $t+1$ the policy arrived closer to accomplishing the desired goal. Before starting Stage 2, we make a copy of the Stage 1 model checkpoint and keep it frozen to use as the reward model for RL training. Environment Definition To implement video generation as environment transitions, we expose the inference interface of the video generation model through remote procedure call, and use the DeepMind RL Environment API (also known as DM Env API) [77] to wrap the remote procedure call in the step function of the environment. When the environment is reset to start a new episode, a goal instruction is randomly sampled from the ones available in the dataset of demonstrations used in Stage 1. RL Method We initialize the RL trained policy using the Stage 1 checkpoint, which as mentioned was also trained with a Behavioral Cloning objective. A collection of actor processes perform policy rollouts in the video generation environment, and add rewards to the trajectories using the reward model defined above. The policy is updated using the REINFORCE [39] objective, i.e. $\nabla_{\pi}\mathcal{L}(o_{t},a_{t},g)=\nabla_{\pi}\log\pi(a_{t}|o_{t},g)\cdot\left[\sum_{i=t}^{T}\gamma^{i-t}\cdot r(o_{i},a_{i},o_{i+1},g)\right]$, where $\mathcal{L}(o_{t},a_{t},g)$ represents the loss associated with the observation-action pair $(o_{t},a_{t})$ in an episode with the goal $g$. The actors are rate limited to prevent generated trajectories from being very off-policy. We report the hyperparameters associated with RL training in Table 7.
@@
11.3 Details of Video Captioning
Note that even though UniSim is a video based simulator trained to condition on past history, we can achieve text-only conditioning by inputting placeholder frames such as white images while increasing the classifier-free guidance strength on text. We found this to work well in generating videos purely from captions of ActivityNet Captions. For generating data to train VLMs, we take the training split of ActivityNet Captions which consists of 30,740 text-video examples after the 50/25/25% train/val1/val2 split as in chen2023pali [40]. For each of the 30,740 text, we generate 4 videos from UniSim, and use the text labels as supervision in finetuning PaLI-X. As a result, we have 4X amount of the original training data (in terms the number of videos). In addition, we found the generated videos to generally align better semantically than the original ActivityNet Captions videos, which could contain noise and ambiguous videos that could be labeled differently. We use ground truth temporal proposals at evaluation following chen2023pali [40] and krishna2017dense [7]. Following chen2023pali [40] and wang2021end [78], we use the val1 split for validation and val2 split for testing.
12 Additional Ablations
12.1 Ablations of Datasets
@@
UniSim without internet 307.80 21.99
UniSim 211.30 22.63
@@
12.2 Ablations of Model Size
@@
13 Failed Simulations without Joint Training
@@
Figure 14: When the text-to-video model behind UniSim is only trained on data from brohan2022rt [19] as opposed incorporating broad data from the internet and other manipulation datasets, long-horizon interaction simulations fail half of the time (red text).

### v3

@@ D.2 Details of RL Policy Training
Stage 1 (Supervised Learning) Model Architecture The PaLI 3B model trained on Language-Table uses a Vision Transformer architecture G/14 (Zhai et al., 2022) to process images, and the encoder-decoder architecture of UL2 language model (Tay et al., 2022) for encoding task descriptions and decoding tokens which can represent language, control actions, or other values of interest (described below). Objectives In the first stage of training, using a dataset of demonstrations, we finetune the pretrained PaLI 3B vision language model checkpoint (Chen et al., 2022b) with the following tasks:
@@
  • Behavioral Cloning: Given observations and task instruction, predict the demonstration action. The continuous actions of the Language-Table domain are discretized into the form “+1 -5”, and represented using extra tokens from the PaLI model’s token vocabulary. As an example, “+1 -5” is represented by the token sequence (<extra_id_65>, <extra_id_1>, <extra_id_66>, <extra_id_5>).
@@
We use learning rate 0.001, dropout rate 0.1, and batch size 128 to finetune the PaLI 3B model for 300k gradient steps with 1k warmup steps on both the simulated and real Language Table dataset similar to RT-2 Brohan et al. (2023).
Stage 2 (RL Training) Reward Definition As mentioned above, during Stage 1, given an observation and goal, the PaLI model is finetuned to predict how many timesteps are left until the demonstration episode reaches a success state. Let us denote this function by $d(o,g)$. The reward we use during RL training is defined as $r(o_{t},a_{t},o_{t+1},g)=-[d(o_{t+1},g)-d(o_{t},g)]\cdot\mathcal{C}$, where $\mathcal{C}>0$ is a small constant used to stabilize training ($\mathcal{C}=5e-2$ in this work). Intuitively, this reward tracks if from timestep $t$ to $t+1$ the policy arrived closer to accomplishing the desired goal. Before starting Stage 2, we make a copy of the Stage 1 model checkpoint and keep it frozen to use as the reward model for RL training. Environment Definition To implement video generation as environment transitions, we expose the inference interface of the video generation model through remote procedure call, and use the DeepMind RL Environment API (also known as DM Env API) (Tassa et al., 2018) to wrap the remote procedure call in the step function of the environment. When the environment is reset to start a new episode, a goal instruction is randomly sampled from the ones available in the dataset of demonstrations used in Stage 1. RL Method We initialize the RL trained policy using the Stage 1 checkpoint, which as mentioned was also trained with a Behavioral Cloning objective. A collection of actor processes perform policy rollouts in the video generation environment, and add rewards to the trajectories using the reward model defined above. The policy is updated using the REINFORCE (Williams, 1992) objective, i.e. $\nabla_{\pi}\mathcal{L}(o_{t},a_{t},g)=\nabla_{\pi}\log\pi(a_{t}|o_{t},g)\cdot\left[\sum_{i=t}^{T}\gamma^{i-t}\cdot r(o_{i},a_{i},o_{i+1},g)\right]$, where $\mathcal{L}(o_{t},a_{t},g)$ represents the loss associated with the observation-action pair $(o_{t},a_{t})$ in an episode with the goal $g$. The actors are rate limited to prevent generated trajectories from being very off-policy. We report the hyperparameters associated with RL training in Table 7.
@@
D.3 Details of Video Captioning
Note that even though UniSim is a video based simulator trained to condition on past history, we can achieve text-only conditioning by inputting placeholder frames such as white images while increasing the classifier-free guidance strength on text. We found this to work well in generating videos purely from captions of ActivityNet Captions. For generating data to train VLMs, we take the training split of ActivityNet Captions which consists of 30,740 text-video examples after the 50/25/25% train/val1/val2 split as in Chen et al. (2023). For each of the 30,740 text, we generate 4 videos from UniSim, and use the text labels as supervision in finetuning PaLI-X. As a result, we have 4X amount of the original training data (in terms the number of videos). In addition, we found the generated videos to generally align better semantically than the original ActivityNet Captions videos, which could contain noise and ambiguous videos that could be labeled differently. We use ground truth temporal proposals at evaluation following Chen et al. (2023) and Krishna et al. (2017). Following Chen et al. (2023) and Wang et al. (2021), we use the val1 split for validation and val2 split for testing.
Appendix E Additional Ablations
E.1 Ablations of Datasets
@@
Without internet 307.80 21.99
Universal simulator 211.30 22.63
@@
E.2 Ablations of Model Size
@@
Appendix F Failed Simulations without Joint Training
@@
Figure 14: When the text-to-video model behind UniSim is only trained on data from Brohan et al. (2022) as opposed incorporating broad data from the internet and other manipulation datasets, long-horizon interaction simulations fail half of the time (red text).